OpenStudy (anonymous):

A ladder leaning against a wall makes a 60o angle with the ground. The base of the ladder is 4 m from the building. How high above the ground is the top of the ladder? 4. Use the Pythagorean Theorem

OpenStudy (anonymous):

so you are going to use trig to solve this problem. using cos=adj/hyp you can find the length of the ladder (hyp). Then once you have that you can use good ol pythag and do a^2+b^2=c^2. The variable a can be the 4m and c is the length of the ladder you just found